The Henry's law constant for CO2 is 3.4 × 10−2 M/atm at 25 °C. Assuming ideal solution behavior, what pressure of carbon dioxide is needed to maintain a CO2 concentration of 0.10 M in a can of lemon-lime soda?
1
Expert's answer
2020-07-14T15:17:23-0400

According to Henry's Law:

P = KH* C

C = concentration

P = 3.4*10-2 *0.1 atm

P = 3.4* 10-3 atm